Nov 3, 2016 Calorimeter

From GlueXWiki
Jump to: navigation, search

Video Conferencing Information

Meeting Time: 11:00 a.m.

  1. To join via a Web Browser, go to the page [1]
  2. To join via Polycom room system go to the IP Address: ( and enter the meeting ID: 907185247.
  3. To join via phone, use one of the following numbers and the Conference ID: 907185247.
    • US or Canada: +1 408 740 7256 or
    • US or Canada: +1 888 240 2560
  4. Upon connection all microphones are automatically muted. To unmute your mike on a Polycom or equivalent unit, enter *4. Unmuting on a computer is trivial as there is a microphone button than can be clicked.
  5. More information on connecting to bluejeans is available.

Participant Direct Lines

  • JLab Phone: in CC F326 is 757-269-6460 (usual room)
  • JLab Phone in CC L207 is 757-269-7084
  • Phone in the Regina Video-conference Suite is 306-585-4204
  • Athens Phone: in Christina's office is 011-30-210-727-6947

Action Items

  • Produce documentation for implementation of timing algorithms and constants (Andrei)
  • Report back on algorithm for track matching to BCAL (Sean)
  • Propose method to track the LED monitoring scheme on a routine basis (Ahmed and Zisis)


  1. FCAL HDFCAL log book
  2. BCAL HDBCAL log book
  3. BCAL Action Items and Working Tasks

Tentative Agenda

  1. Announcements
  2. Action Items
  3. FCAL update
  4. BCAL update
  5. Calibrations
  6. Reconstruction and Simulation
  7. Any other business


Attendees: Elton, Mark, Jon, Adesh, Mike (JLab); Christina, George (Athens); Will (CMU); Zisis, Ahmed, Andrei (UofR); Sean (NW); Matt (IU).

  1. Announcements
    • Accelerator is making good strides toward fixing the power supply in arc 7. They expect to start earlier than originally expected. Alexandre has sent out appropriate announcements.
  2. Action Items
    • Sean: Track matching to downstream end of BCAL was an issue. The new tracking is much better, but matching to the downstream end of BCAL needs checking with the latest launch.
  3. FCAL update
    1. Bases (Jon)
      • Replaced 36 bases this week with Chris.
      • This was productive training for Chris since most variations of replacement had to be addressed.
      • Have about 30 remaining spares
      • Chris has repaired 15 bases to date. One has failed, but hopefully this was due to poor implementation and not the procedure itself.
      • Meeting (Adesh, Jon, Elton, Paul Smith, Chris) with Fernando and Matt tomorrow to go over procedures and make sure everyone is on the same page.
    2. Preparations (Adesh)
      • DAQ and pulser are operational
      • Scalers (EPICs) were not working
      • Will work on the Hot checkout today
  4. BCAL update (Mark and Elton)
    1. Discriminator scans (Mark)
      • Discriminator threshold scans of baseline were completed on all channels. Found some baseline above zero.
      • In order to optimize thresholds, require flexibility for negative threshold settings. These have been implemented and need checking
      • Most recent cosmic-ray data show very uneven response in the TDC occupancy plots. Need to recheck baseline scans
    2. Occupancies
      • The fadc occupancy seemed uneven on the upstream, more uniform on the downstream.
      • Could be due to pedestal settings, which Serguei will update shortly.
      • Scalers also need to be checked as they were not all working.
    3. New integration windows:
      • Settings for new integration window will be 2-13.
      • Zisis: Should the cosmic data be analyzed for calibration with muons?
      • Will: The pi0 peak selection can now fairly easy start from previous settings, even if the scale is changed.
      • The plan is to start with scaled constants from the old to new integration windows. Scale will be estimate in the next couple of days.
      • Constants will be determined starting from 2 different initial values.
    4. Monitoring (Ahmed)
      • Latest data has not yet been checked
      • Need to upload plugin and display scripts to sim-recon / git.
      • Reference values will be stored initially on the work disk and monitoring will be played offline.
      • Sean will review code once it is checked into git and consider the best option for storing references in CCDB using JANA resources.
  5. Reconstruction and Simulation
    • BCAL Smearing update (Will M.)
    • Removed unused simulation constants in the smearing code
    • Moved all active constants to CCDB.
    • Some questions: BCAL_mevPerPE unknown origin -> Andrei will provide reference to Will
    • Updated simulated resolutions generated with reasonable parameters now track data for pi0 and design for deltaE/E.
    • Constants still need to be fine tuned to have MC match data.
    • Zisis: Could you extend calculation to higher energy? Will: yes
  6. Calibrations
    • Studies of BCAL timing with ADCs and TDCs (Mark)
    • Mark reviewed his slides, here are a few comments and questions
    • Andrei: Why not fit time-walk to lower ADC values? Mark: The system has not control of the baseline and with a constant threshold, channel to channel variations are large. Fit only over reliable range for TDC. With the new run we expect to be able to adjust thresholds individually and we will have better uniformity. For very small ADC values, the time is taken from the ADCs. Also use the ADC time if there is a large discrepancy between TDC and ADC times.
    • Shower timing: Present algorithm is to extrapolate all times to the inner radius of the BCAL and then average using E^2 weighting.
    • Andrei: Suggest E weight for average may be better. Mark: Will try various options
    • DeltaT vs energy: charged DeltaT decreases and then increases with energy (not understood). Neutral DeltaT decreases with energy, as expected
    • Point time offsets: use position of point instead of position of shower
    • Conclusion: Not yet achieving full resolution out of the BCAL
    • Will: Correction to curvature inside BCAL (did not do this a couple of years ago)? Mike: Possible that tracks follow magnetic field up to the BCAL only. Need to check with Simon.
    • Elton asks Andrei: Does your shower time increase at high energy? Andrei: No. Timing resolution follows ~ 100 ps/sqrt(E) + c and decreases with energy.
    • Andrei: Asked about the systematics of charged track time resolution with z. His algorithm has smaller systematics increasing from about 200 ps to 350 ps near the downstream of BCAL.
    • Andrei: Did George have to redo his fits for veff after Mark's adjustment of ADC offsets? Mark: No, since the two sets of constants are orthogonal to each other.
    • Andrei: Do you use single ended hits? Mark: If there is no TDC time on one side, the ADC time is used. If there his no ADC or TDC, the hit is not used in the shower timing
    • Andrei: Adjustment of ADC offsets by Mark significantly improved ADC timing and brings them closer to the TDC time resolutions
  7. Any other business